OpenStudy (anonymous):

Write the equation of a rational function that has vertical asymptote at x=-5 and x=0, a horizontal asymptote at y=4, x-intercepts at (-1,0) and (1,0), and no y-intercept. Please help!

OpenStudy (tkhunny):

Rational Function... vertical asymptote at x=-5 ==> \(y = \dfrac{1}{x+5}\) vertical asymptote at x=0 ==> \(y = \dfrac{1}{x(x+5)}\) Horizontal at y = 4 ==> \(y = \dfrac{4(x-2)^{2}}{x(x+5)}\) <== This was an arbitrary choice. The only WRONG numerator would be a factor of 'x' or 'x+5' x-interecept at x = 1 ==> \(y = \dfrac{4(x-2)(x-1)}{x(x+5)}\) x-interecept at x = -1 ==> \(y = \dfrac{4(x+1)(x-1)}{x(x+5)}\) Just one piece at a time!

OpenStudy (anonymous):

Thank you!

OpenStudy (tkhunny):

That was your one free one. Let's see you do one.